OpenStudy (anonymous):

The student council wants to buy some stuffed animals to give to children at the local hospital. Stuffed bears cost $6 each and stuffed dogs each cost $4 from a wholesale catalog. The student council wants to spend no more than $170 but at least $80 on the toys. a. Let x represent the number of bears and y represent the number of dogs. Write two inequalities that represent the two amounts the council wants to spend. Explain what each part of the inequalities means. Find one solution for this problem. Explain how you found the solution. What is the total cost for this solution?

sam (.sam.):

the council doesnt want to spend more than 170$ on the toys. \[6x + 4y \le 170\] The council also wants to spend at least 80$ on the toys. \[6x + 4y \ge 80\] 6x = the number of bears that will be bought multiplied by the price of the bear 4y= the number of dogs that will be bought multiplied by the price of the dog then let x = 10 and y = 10 (meaning you buy 10 bears and dogs) You will spend a total of 6(10) + 4(10) 60 + 40 = 100 100 is less than 170, but greater than 80, so its valid solution

OpenStudy (anonymous):

Thank you!

sam (.sam.):

yw